OpenStudy (anonymous):

Explain how estimation could be used to make sure the answer is reasonable. Gord mixed 2.45 L of orange juice with 3.65 L of pineapple juice. His friends drank 4.85 L of the juice mixture. How much juice was left over?

OpenStudy (anonymous):

A. Answer: The amount left over was 1.25 L. Estimation: 2.45 + 3.65 is about 6. 6 L – 5 L = 1 L There was a bit more than 1 L left over. B. Answer: The amount left over was 2.75 L. Estimation: 2.45 + 3.65 is about 7. 7 L – 5 L = 2 L There was a bit more than 2 L left over. C. Answer: The amount left over was 10.95 L. Estimation: 2.45 + 3.65 is about 6 L. 6 L + 5 L = 11 L There was a bit less than 11 L left over.

OpenStudy (anonymous):

2.45+3.65

OpenStudy (anonymous):

6.1

OpenStudy (anonymous):

which is about what?

OpenStudy (anonymous):

would you round up or down?

OpenStudy (anonymous):

down, 6

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

would you round up or down 4.85?

OpenStudy (anonymous):

be right back

OpenStudy (anonymous):

yes, to 5

OpenStudy (anonymous):

ok, so what would be the answer?

OpenStudy (anonymous):

b?

OpenStudy (anonymous):

If you start with 6 and take away 5, how many do you have left?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

so, A?

OpenStudy (anonymous):

that's right, good job!

OpenStudy (anonymous):

Thank you for the help @10115658

OpenStudy (anonymous):

no problem